OpenStudy (anonymous):

x/8 + 3= x/8 Need help in solving please

OpenStudy (anonymous):

x=1/0 :)

OpenStudy (anonymous):

this question doesn't ake any sense!!!

OpenStudy (anonymous):

pls check ur ques

OpenStudy (anonymous):

\[x/8 + 3= x/8\]

OpenStudy (anonymous):

@thedane listen carefully let x/8=t then you want to say if we add 3 to a no. t it doesn't affect it so either the question is wrong or it means that the no. is so large that it remains unaffected so it would be infinity that means 1/0:)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome:)